The formula C=2pir is used to calculate the circumference of a circle when given the radius. What is the formula for calculating the radius when give the circumference?
A. C/2pi =r
B. C/2r =pi
C. 2piC=r
D. 2pi/C=r

Answer by stanbon(75887)

The formula C=2pir is used to calculate the circumference of a circle when given the radius. What is the formula for calculating the radius when give the circumference?
C = 2*pi*r
------
Divide both sides by 2pi to get::
r = C/(2pi)
--------------------------
Cheers,
Stan H.
---------------
A. C/2pi =r
B. C/2r =pi
C. 2piC=r
D. 2pi/C=r